[figure: 9.7 cm]

ON THE APPEARANCE OF THE FERROMAGNETIC SHORT-RANGE COUPLING IN THE NEW

ANTIFERROMAGNETIC SEMICONDUCTOR CuCr_{1.6}Sb_{0.4}S_{4}

J. Warczewski^{a}, J. Krok-Kowalski^{a}, L. I. Koroleva^{b}, A. G. Odintsov^{b}, T. Mydlarz^{c}, A. Gilewski^{c} aUniversity of Silesia, Institute of Physics, ul. Uniwersytecka 4, 40–007 Katowice, Poland bM.V. Lomonosov Moscow State University, 119899 Moscow, Russia cInternational Laboratory of High Magnetic Fields and Low Temperatures, ul. Gajowicka 95, 53 529 Wrocław, Poland

It turns out that the compound under study has antiferromagnetic long-range ordering with the Néel temperature equal to about 40 K. One observes also high positive Curie-Weiss paramagnetic temperature Θ_{C−W} equal to 230 K. This value points to the appearance of the strong ferromagnetic short-range coupling. One can also observe on the magnetization curves a hysteresis which appears in the region of magnetic fields up to 3 Tesla. In spite of the existence of Cu ions in the tetrahedral positions a semiconductive character of this compound seems to be connected with the blockade of the double exchange magnetic interaction, the latter effect being caused by Sb ions in the octahedral positions.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Prof. Jerzy Warczewski

Address for correspondence : University of Silesia, Institute of Physics, ul. Uniwersytecka 4, 40007 Katowice, Poland

Email address : warcz@us.edu.pl